OpenStudy (anonymous):

put in slope intercept form: slope 3/2, y-intercept 4

OpenStudy (anonymous):

y=3/2x+4

OpenStudy (anonymous):

the eqn is y = (slope)x + (y-intercept) => y = (3/2)x + 4

OpenStudy (anonymous):

how did you get 4 as b?

OpenStudy (anonymous):

y=3/2 x +4

OpenStudy (anonymous):

y=mx+b. b is the y intercept

OpenStudy (anonymous):

plugging slope in m and y-intercept in b for : y=mx+b